how many different rectangles are there with an area of 36cm squared if the length of each, in centimeters are whole numbers
Answer by Theo(13342)

1 * 36 = 36
2 * 18 = 36
3 * 12 = 36
4 * 9 = 36
6 * 6 = 36

I think that's it.